What is the approximate residual error (e) at C = 100 for the relationship between Celcius (C) and Fahrenheit (F) temperature, given that the formula for the relationship is: F = 32 + (1.8 * C)?
Answer #1

The estimated Fahrenheit (F) temperature = 32 + (1.8*100) = 212

It is known that,

At C = 100, Fahrenheit (F) temperature 212

The approximate residual error (e) at C = 100

= observed - predicted

= 212 - 212

= 0